OpenStudy (rayman5766):

What value of d makes the statement true? d + 12.6 = 9 A. –21.6 B. –3.6 C. 3.6 D. 21.6

OpenStudy (rayman5766):

Sam had a full 1 L bottle of olive oil. He poured 0.25 L of the olive oil into a container to make salad dressing. He also poured an unknown amount x onto a plate as a bread dip. There was 0.6 L of olive oil remaining in the bottle. Which equation can be used to solve for x? A. 1 + 0.25 – x = 0.6 B. 1 + 0.25 + x = 0.6 C. 1 – 0.25 – x = 0.6 D. 1 – 0.25 + x = 0.6

OpenStudy (anonymous):

D+12.6 = 9 ... Subtract 12.6 - 9 ANSWER: C. 3.6